# Hawking radiation in a charged, cosmological black hole

### v1

@@ III.4 Effective temperature perceived by a radial freefaller with negative specific energy
An observer inside a trapping horizon, who continuously accelerates outwards, against the curvature of spacetime acquires a negative energy (state with $\frac{dt}{d\tau}<0$). It is not possible to be outfalling inside a black hole, and therefore, the observer will still be travelling to a smaller spatial radius. But the outward acceleration will cause the observer to pass through the future Cauchy horizon, in contrast to an observer of positive energy, who will travel to the past Cauchy horizon. The outward acceleration by the observer, or the application of an external force may be seen as agents to Lorentz-boost the observer into a frame of negative energy.
@@
The effective temperatures perceived by an outgoing observer within the event horizon is shown in Fig. 3. Such an observer will not see a very significant change in the Hawking modes, except for an interchange in the ingoing and outgoing effective temperatures. As a consequence, it will be the ingoing modes diverging to infinity at the inner horizon and the outgoing modes tending to a finite negative value. The apparent infinite blueshift in the ingoing modes may be attributed to an infinite Doppler shift as the horizon function goes to zero. But it is not only at the event horizon that this observer will see an infinite wall of energy—the horizon function similarly goes to zero at the event horizon, making it impossible for the observer to cross. This can be seen, once again, in the divergence of the ingoing modes at this region. The observer is therefore confined to the region between the event and the inner horizons, consistent with the nature of a trapping horizon. An observer of zero energy is similarly unable to traverse beyond the event horizon. This situation is similar to the charged Nariai limit (in that the observer has zero energy), and indeed, there is thermal equilibrium for the most part, broken only as the observer approaches the event horizon. The cold regime also displays a divergence in the ingoing modes, with the outgoing modes starting to diverge as the energy becomes more negative. At any point inside the black hole, $\kappa_{o}^{-}\propto-E$ for $E<0$ and $k_{i}^{+}\propto E$ for $E>0$. Both $\kappa_{i}^{-}$ and $\kappa_{o}^{+}$ asymptotically tend to zero as $E\xrightarrow{}+\infty$ and $E\xrightarrow{}-\infty$ respectively, as in Figure 4. Notably, the change in energy does not affect either the transition points or the sign in either of the modes. Neither the divergence at the inner horizon, nor the Hawking radiation itself can ever be forced to go to zero by a change in state of the observer.
@@
In an RNdS black hole, there exists another trapping region outside the cosmological horizon. While this paper, in general, does not study this region in great detail, it is worth considering the implications of negative energy states in this region. The observer will be forced to move to a greater spatial radius (outfalling), but can be ingoing. The equilibrium in the modes as the observer tends to infinity remains unchanged, but the outgoing modes diverge as the observer approaches the cosmological horizon, while the ingoing modes tend to zero. Once again, this is consistent with the expectations that the observer should not be able to cross the cosmological horizon back into the observable universe. Unlike inside the event horizon, the effective temperatures in this region are independent of the specific energy in the infinite limit. However, the convergent distance for the ingoing and outgoing modes becomes farther with increasingly negative energy, and the divergence gets sharper towards the cosmological horizon. But once again this implies that it is not possible to Lorentz-boost an observer into a vacuum state that persists over all of spacetime for any parametric values of the black hole.
@@ IV Effective temperature perceived by radial free fallers in RNAdS spacetime($\Lambda<0$)
Having discussed, at great length, the effective temperatures perceived by an observer in de Sitter space, we now turn our attention to the Anti-de Sitter (AdS) space. AdS space has been long regarded as highly unrealistic, two reasons for which were cited by Hawking and Page [51]: first, a negative cosmological constant corresponds to a negative energy density; second, closed timelike curves were seen to form on a manifold in the AdS space. While the second reason provided has been since invalidated [52], solving the apparent paradox by unwrapping the time-like circles and mapping them onto the real numbers ($\infty$- ∞ to +$\infty$+ ∞), the first reason still appears to stand. However, recent developments in quantum gravity proposed the holographic principle, a potential resolution for the information paradox. It involves a dimensional reduction to encode information from the bulk of an N dimensional space onto the boundary of the space [53]. Interest in the AdS space stems from the ability of the AdS/CFT correspondence to form a map from the bulk to the boundary, enabling the dimensional reduction [54, 55].
@@ IV.1 Vacuum State
A major distinction between the setup for the RNdS metric and the RNAdS metric comes from the fact that the vacuum state cannot be defined as easily in a contracting universe as it is in an expanding one. The Unruh vacuum, as described earlier, requires that the outgoing modes be positive with respect to the past horizon, and the ingoing modes be positive frequency with respect to $\partial/\partial t$ at past null infinity which would generally be the outer boundary of the spacetime. But this boundary is infinity, which gets compressed as the space contracts. Since $\Delta\xrightarrow{}\infty$ as $r\xrightarrow{}\infty$, a clock on any emitter placed at this boundary will have to tick infinitely faster than the global time coordinate. So while the Unruh vacuum can be defined with respect to the past horizon, it is ill defined at the AdS boundary. The eternal AdS geometry therefore cannot be modelled in entirety by the Unruh vacuum.
@@
A very similar proof as in Section II.3 can be used to show that the proper time of an emitter in free fall at the event horizon is proportional to the outgoing Kruskal coordinate at the event horizon. The outgoing modes will travel out to the boundary of the AdS spacetime at infinity, from where they will be reflected back as ingoing modes. The AdS spacetime may be viewed as a finite box, which will slowly get filled with radiation [56]. At some point in process, a thermal equilibrium is reached, where the ingoing and outgoing modes have the same effective temperature. The configuration is modelled best by the Hawking-Hartle state, where the modes are positive with respect to the affine parameters on their respective boundaries of generation. [57] offers a more detailed insight into the mode solutions in AdS spacetime.
@@ IV.2 Effective Temperature
$\kappa_{i}^{+}=-\frac{E+\sqrt{E^{2}-\Delta}}{\Delta L_{AdS}}\left(\varkappa_{ob}-\varkappa_{+}\right),$ (37a)
$\kappa_{i}^{in}=\frac{E-\sqrt{E^{2}-\Delta}}{\Delta L_{AdS}}\left(\varkappa_{ob}-\varkappa_{+}\right).$ (37b)
A change in notation from superscript c to “in” has been introduced to reinforce the fact that the ingoing modes are no longer generated from the cosmological horizon. The only difference between these equations and the ones for a dS spacetime is that the equation for the ingoing modes depends upon $\varkappa_{+}$ instead of $\varkappa_{c}$. The redshift experienced by the modes on the way out from the position of the observer to infinity is cancelled by a corresponding blueshift as it reflects and travels back. The net red/blueshift observed is only that which occurs between the event horizon and the position of the observer. It is akin to the ingoing modes having been generated from the event horizon itself. The equations implicitly assume that the outgoing modes are radial in the spherical geometry, and undergo perfect reflection at the AdS boundary with a zero phase shift.
@@
The simplest case of an RNAdS black hole, is once again, when the charge goes to zero, which falls under Schwarzschild-Anti-de Sitter (SAdS) limit, the graph of which is on the upper left of Figure 5. Having analyzed in Section III, the effects of the mass and positive cosmological constant, the SAdS limit provides the perfect context in which to examine the effects of the cosmological contraction, in isolation from the effects of mass and charge. Knowing that the mass causes a redshift, the negative effective temperatures in the outgoing modes may be attributed entirely to the effects of the negative cosmological constant. The same assumption cannot be made for the ingoing modes, given that their source of origin is not the same as it was in the SdS case. The effects of this will be analyzed separately. As the observer starts out from rest and falls inwards, travelling in predominantly AdS space, the outgoing modes appear blueshifted, with the rate of blueshift decreasing as the effects of mass takeover. It is not unexpected that if $\Lambda>0$ results in a redshift, that $\Lambda<0$ causes a blueshift. The outgoing modes are decreasingly redshifted away from the event horizon, and become increasingly blueshifted as they travel out to infinity. Having reflected off the boundary, they then become ingoing modes. At this point, the frequency of the modes are traversing a slope of the horizon function greater than that at the event horizon, resulting in a gravitational redshift. In addition, the observer is also ingoing, resulting in a Doppler redshift as well. Effectively, it is the source of the ingoing modes that results in the perceived redshift in AdS space. As they near the event horizon, they start to experience a gravitational blueshift. Like in the previous cases, the point of transition from positive to negative effective temperatures occurs where $\varkappa_{ob}=\varkappa_{+}$. Because both the ingoing and outgoing modes source from $r_{+}$, they both undergo transitions at exactly the same points, except for the transition of the ingoing modes at the event horizon (a consequence of a zero Doppler shift). The mass predominates beyond the event horizon, and the graph, in this regime, bears close resemblance to those of the SdS and Schwarzschild cases [11], where the effective temperatures are positive and eventually converge, with the sharpness of the convergence being proportional to M. There are also no parameters at which the SAdS black hole exists in thermal equilibrium, and except for at the rest point, the observer never perceives an equal effective temperature (magnitude) in the ingoing and outgoing modes. This is once again a result of the family of emitters for the ingoing modes (the Doppler shift factors are never equal, but the gravitational redshift factors are always equal).
@@
The upper right graph of Figure 5 represents a fairly general case of the RNAdS parameter space. The properties of the inner horizon are the same as in the RNdS metric, i.e. it will have a negative surface gravity, thereby repelling the infalling modes. The divergence of the outgoing modes at the horizon, and the blueshift in the ingoing modes, therefore remains the same. While the general behaviour of the modes up to the blueshift at the inner horizon does not change, the relative distances of the transition points change depending on the cosmological constant and the $Q/M$ ratio. As the cosmological contraction increases ($\Lambda$ decreases by becoming more negative, and the scaled mass parameter $M$ increases), the outer transition point comes closer to the event horizon as would be expected if the AdS nature of spacetime takes precedence over mass closer to the black hole. As the inner and event horizons approach each other, it is possible for the ingoing modes to continue as a negative value across the event horizon, approaching the degenerate limit of M= Q as in lower graph of the figure. This critical ratio for this continuity decreases as the cosmological constant gets more negative. The extent of the intermediate redshift (between $r_{-}$ and $r_{+}$) is seen to decrease as the M/Q ratio decreases, once again consistent with the fact that a larger charge parameter should contribute to a greater blueshift.
@@ V Validity of the adiabatic approximation
The overdot represents the derivative with respect to the proper time of the observer/emitter. Neither the analytic forms of the control function, nor its limit under any special circumstances are straightforward enough to be directly enlightening. However, the general behaviour of $\epsilon$ may be inferred from the graphs in Figures 6 and 7. The superscript convention for the control function $\epsilon$ will follow that introduced for the corresponding effective temperature functions in both the RNdS and RNAdS models.
@@
Adiabaticity is maintained to the greatest extent for an outfaller beyond the rest point. $\epsilon$ is continuous across the cosmological horizon and into the limit of a predominantly de Sitter space. Generally, the control function goes to a minimum at the rest point of the observer ($\dot{r}_{ob}=0$). The exception is in the extremal case (right panel in Figure 6), where the control function diverges for outgoing modes, due to the transition in the effective temperature of the outgoing modes occurring exactly at the degenerate horizons ($r_{0}=r_{+}=r_{c}$). $\epsilon$ diverges as the effective temperatures go to zero; the formalism is therefore least valid near the transition points. For general RNdS models, both $\epsilon^{+}$ and $\epsilon^{c}$ are close to the zeroth order of magnitude inside the event horizon and rise sharply at the transition points before coming back down the zeroth order close the Cauchy horizon. But the zeroth order of magnitude is still greater than one, indicating that adibaticity is not maintained. There are therefore, only three regions in which the approximation holds: at and beyond the rest point, slightly before the transition points where $\frac{d\kappa}{dr_{ob}}\approx 0$ and close to the inner horizon. In the SdS limit, however, there being no inner horizon, $\lim_{r\to 0}(\epsilon^{+},\epsilon^{c})=3$. The control function is only less than one in the proximity of the rest point.
@@
In Anti-de Sitter space, the ingoing modes are, in general, adiabatic in only two regimes: first is near the event horizon, at the minima of the effective temperature function, second is at the local maxima before the inner horizon. For an SAdS model, both $\epsilon^{+}$ and $\epsilon^{in}$ converge as they approach the singularity. The asymptotic limit is proportional to the mass parameter, and is less than 0.5 for $M<2$. As the mass parameter increases, the cosmological constant becomes less negative (i.e. increases), and the contraction reduces. As the contraction reduces, adibaticity is maintained to a greater extent inside the event horizon. It also generally appears to be true that the outgoing modes are more adiabatic than the ingoing modes. This is consistent with the graphs from Figure 5, according to which the magnitude of the outgoing modes is greater than the ingoing modes.
@@
For the most part of the observer’s trajectory, the adiabatic condition does not hold. One may question the implications for the interpretation of $\kappa$. The form of $\kappa$, as in Eq. (23) and used throughout the paper, is derived directly from the definition given by Eq. (1). It does not rely upon the adiabatic approximation. The observer will still see a nonzero particle production, and therein the Hawking radiation. But the radiation will not have a thermal blackbody spectrum. The graphs in Figures 6 and 7 show the regime in which the radiation is thermal.

### v2

@@ III.4 Effective temperature perceived by a radial freefaller with negative specific energy
An observer inside a trapping horizon, who continuously accelerates outward, against the curvature of spacetime acquires a negative energy (state with $\frac{dt}{d\tau}<0$). It is not possible to be outfalling inside a black hole, and therefore, the observer will still be traveling to a smaller spatial radius. But the outward acceleration will cause the observer to pass through the future Cauchy horizon, in contrast to an observer of positive energy, who will travel to the past Cauchy horizon. The outward acceleration by the observer, or the application of an external force may be seen as agents to Lorentz-boost the observer into a frame of negative energy.
@@
Figure 6: A Penrose diagram representing the trajectory of an observer with a negative specific energy beyond the event horizon. As in the diagram, the observer will eventually reach the right section of the inner horizon.
@@
The effective temperatures perceived by an outgoing observer within the event horizon is shown in Fig. 4, and the Penrose diagram corresponding to the situation is given in Fig. 6. Such an observer will not see a very significant change in the Hawking modes, except for an interchange in the ingoing and outgoing effective temperatures. As a consequence, it will be the ingoing modes diverging to infinity at the inner horizon and the outgoing modes tending to a finite negative value. The apparent infinite blueshift in the ingoing modes may be attributed to an infinite Doppler shift as the horizon function goes to zero. But it is not only at the event horizon that this observer will see an infinite wall of energy—the horizon function similarly goes to zero at the event horizon, making it impossible for the observer to cross. This can be seen, once again, in the divergence of the ingoing modes at this region. The observer is therefore confined to the region between the event and the inner horizons, consistent with the nature of a trapping horizon. An observer of zero energy is similarly unable to traverse beyond the event horizon. This situation is similar to the charged Nariai limit (in that the observer has zero energy), and indeed, there is thermal equilibrium for the most part, broken only as the observer approaches the event horizon. The cold regime also displays a divergence in the ingoing modes, with the outgoing modes starting to diverge as the energy becomes more negative. At any point inside the black hole, $\kappa_{o}^{-}\propto-E$ for $E<0$ and $k_{i}^{+}\propto E$ for $E>0$. Both $\kappa_{i}^{-}$ and $\kappa_{o}^{+}$ asymptotically tend to zero as $E\xrightarrow{}+\infty$ and $E\xrightarrow{}-\infty$ respectively, as in Fig. 5. Notably, the change in energy does not affect either the transition points or the sign in either of the modes. Neither the divergence at the inner horizon, nor the Hawking radiation itself can ever be forced to go to zero by a change in state of the observer.
@@
In an RNdS black hole, there exists another trapping region outside the cosmological horizon. While this paper, in general, does not study this region in great detail, it is worth considering the implications of negative energy states in this region. The observer will be forced to move to a greater spatial radius (outfalling), but can be ingoing. The equilibrium in the modes as the observer tends to infinity remains unchanged, but the outgoing modes diverge as the observer approaches the cosmological horizon, while the ingoing modes tend to zero. Once again, this is consistent with the expectations that the observer should not be able to cross the cosmological horizon back into the observable universe. Unlike inside the event horizon, the effective temperatures in this region are independent of the specific energy in the infinite limit. However, the convergent distance for the ingoing and outgoing modes becomes farther with increasingly negative energy, and the divergence gets sharper toward the cosmological horizon. But once again this implies that it is not possible to Lorentz-boost an observer into a vacuum state that persists over all of spacetime for any parametric values of the black hole.
@@ IV Effective temperature perceived by radial free fallers in RNAdS spacetime($\Lambda<0$)
Having discussed, at great length, the effective temperatures perceived by an observer in de Sitter space, we now turn our attention to the anti-de Sitter (AdS) space. AdS space has been long regarded as highly unrealistic, two reasons for which were cited by Hawking and Page [56]: first, a negative cosmological constant corresponds to a negative energy density; second, closed timelike curves were seen to form on a manifold in the AdS space. While the second reason provided has been since invalidated [57], solving the apparent paradox by unwrapping the timelike circles and mapping them onto the real numbers ($\infty$- ∞ to +$\infty$+ ∞), the first reason still appears to stand. However, recent developments in quantum gravity proposed the holographic principle, a potential resolution for the information paradox. It involves a dimensional reduction to encode information from the bulk of an N dimensional space onto the boundary of the space [58]. Interest in the AdS space stems from the ability of the AdS/CFT correspondence to form a map from the bulk to the boundary, enabling the dimensional reduction [59, 60].
@@ IV.1 Vacuum state
A major distinction between the setup for the RNdS metric and the RNAdS metric comes from the fact that the vacuum state cannot be defined as easily in a contracting universe as it is in an expanding one. The Unruh vacuum, as described earlier, requires that the outgoing modes be positive with respect to the past horizon, and the ingoing modes be positive frequency with respect to $\partial/\partial t$ at past null infinity which would generally be the outer boundary of the spacetime. But this boundary is infinity, which gets compressed as the space contracts. Since $\Delta\xrightarrow{}\infty$ as $r\xrightarrow{}\infty$, a clock on any emitter placed at this boundary will have to tick infinitely faster than the global time coordinate. So while the Unruh vacuum can be defined with respect to the past horizon, it is ill defined at the AdS boundary. The eternal AdS geometry therefore cannot be modeled in entirety by the Unruh vacuum.
@@
A very similar proof as in Section II.2 can be used to show that the proper time of an emitter in free fall at the event horizon is proportional to the outgoing Kruskal coordinate at the event horizon. The outgoing modes will travel out to the boundary of the AdS spacetime at infinity, from where they will be reflected back as ingoing modes. The AdS spacetime may be viewed as a finite box, which will slowly get filled with radiation [61]. At some point in process, a thermal equilibrium is reached, where the ingoing and outgoing modes have the same effective temperature. The configuration is modeled best by the Hawking-Hartle state, where the modes are positive with respect to the affine parameters on their respective boundaries of generation. [62] offers a more detailed insight into the mode solutions in AdS spacetime.
@@ IV.2 Effective temperature
Figure 7: Both the observer (orange) and the emitter (purple) are ingoing, with the observer having been released from an arbitrary point well outside the event horizon. Outgoing modes are always produced from the ingoing family of emitters, but may reflect off the AdS boundary to yield the ingoing modes.
@@
$\kappa_{i}^{+}(r_{ob})=-\frac{E+\sqrt{E^{2}-\Delta(r_{ob})}}{\Delta(r_{ob})L_{AdS}}\left(\varkappa(r_{ob})-\varkappa_{+}\right),$ (39a)
$\kappa_{i}^{in}(r_{ob})=\frac{E-\sqrt{E^{2}-\Delta(r_{ob})}}{\Delta(r_{ob})L_{AdS}}\left(\varkappa(r_{ob})-\varkappa_{+}\right).$ (39b)
A change in notation from superscript c to “in” has been introduced to reinforce the fact that the ingoing modes are no longer generated from the cosmological horizon. The only difference between these equations and the ones for a dS spacetime is that the equation for the ingoing modes depends upon $\varkappa_{+}$ instead of $\varkappa_{c}$. The redshift experienced by the modes on the way out from the position of the observer to infinity is canceled by a corresponding blueshift as it reflects and travels back, as shown in Fig. 7. The net red/blueshift observed is only that which occurs between the event horizon and the position of the observer. It is akin to the ingoing modes having been generated from the event horizon itself. The equations implicitly assume that the outgoing modes are radial in the spherical geometry, and undergo perfect reflection at the AdS boundary with a zero phase shift.
@@
The simplest case of an RNAdS black hole, is once again, when the charge goes to zero, which falls under Schwarzschild-anti-de Sitter (SAdS) limit, the graph of which is on the upper left of Fig. 8. Having analyzed in Section III, the effects of the mass and positive cosmological constant, the SAdS limit provides the perfect context in which to examine the effects of the cosmological contraction, in isolation from the effects of mass and charge. Knowing that the mass causes a redshift, the negative effective temperatures in the outgoing modes may be attributed entirely to the effects of the negative cosmological constant. The same assumption cannot be made for the ingoing modes, given that their source of origin is not the same as it was in the SdS case. The effects of this will be analyzed separately. As the observer starts out from rest and falls inward, traveling in predominantly AdS space, the outgoing modes appear blueshifted, with the rate of blueshift decreasing as the effects of mass takeover. It is not unexpected that if $\Lambda>0$ results in a redshift, that $\Lambda<0$ causes a blueshift. The outgoing modes are decreasingly redshifted away from the event horizon, and become increasingly blueshifted as they travel out to infinity. Having reflected off the boundary, they then become ingoing modes. At this point, the frequency of the modes are traversing a slope of the horizon function greater than that at the event horizon, resulting in a gravitational redshift. In addition, the observer is also ingoing, resulting in a Doppler redshift as well. Effectively, it is the source of the ingoing modes that results in the perceived redshift in AdS space. As they near the event horizon, they start to experience a gravitational blueshift. Like in the previous cases, the point of transition from positive to negative effective temperatures occurs where $\varkappa(r_{ob})=\varkappa_{+}$. Because both the ingoing and outgoing modes source from $r_{+}$, they both undergo transitions at exactly the same points, except for the transition of the ingoing modes at the event horizon (a consequence of a zero Doppler shift). The mass predominates beyond the event horizon, and the graph, in this regime, bears close resemblance to those of the SdS and Schwarzschild cases [11], where the effective temperatures are positive and eventually converge, with the sharpness of the convergence being proportional to M. There are also no parameters at which the SAdS black hole exists in thermal equilibrium, and except for at the rest point, the observer never perceives an equal effective temperature (magnitude) in the ingoing and outgoing modes. This is once again a result of the family of emitters for the ingoing modes (the Doppler shift factors are never equal, but the gravitational redshift factors are always equal).
@@
The upper right graph of Fig. 8 represents a fairly general case of the RNAdS parameter space. The properties of the inner horizon are the same as in the RNdS metric, i.e. it will have a negative surface gravity, thereby repelling the infalling modes. The divergence of the outgoing modes at the horizon, and the blueshift in the ingoing modes, therefore remains the same. While the general behavior of the modes up to the blueshift at the inner horizon does not change, the relative distances of the transition points change depending on the cosmological constant and the $Q/M$ ratio. As the cosmological contraction increases ($\Lambda$ decreases by becoming more negative, and the scaled mass parameter $M$ increases), the outer transition point comes closer to the event horizon as would be expected if the AdS nature of spacetime takes precedence over mass closer to the black hole. As the inner and event horizons approach each other, it is possible for the ingoing modes to continue as a negative value across the event horizon, approaching the degenerate limit of M= Q as in lower graph of the figure. This critical ratio for this continuity decreases as the cosmological constant gets more negative. The extent of the intermediate redshift (between $r_{-}$ and $r_{+}$) is seen to decrease as the M/Q ratio decreases, once again consistent with the fact that a larger charge parameter should contribute to a greater blueshift.
@@ V Validity of the adiabatic approximation
The overdot represents the derivative with respect to the proper time of the observer/emitter. Neither the analytic forms of the control function, nor its limit under any special circumstances are straightforward enough to be directly enlightening. However, the general behavior of $\epsilon$ may be inferred from the graphs in Figs. 9 and 10. The superscript convention for the control function $\epsilon$ will follow that introduced for the corresponding effective temperature functions in both the RNdS and RNAdS models.
@@
Adiabaticity is maintained to the greatest extent for an outfaller beyond the rest point. $\epsilon$ is continuous across the cosmological horizon and into the limit of a predominantly de Sitter space. Generally, the control function goes to a minimum at the rest point of the observer ($\dot{r}_{ob}=0$). The exception is in the extremal case (right panel in Figure 9), where the control function diverges for outgoing modes, due to the transition in the effective temperature of the outgoing modes occurring exactly at the degenerate horizons ($r_{0}=r_{+}=r_{c}$). $\epsilon$ diverges as the effective temperatures go to zero; the formalism is therefore least valid near the transition points. For general RNdS models, both $\epsilon^{+}$ and $\epsilon^{c}$ are close to the zeroth order of magnitude inside the event horizon and rise sharply at the transition points before coming back down the zeroth order close the Cauchy horizon. But the zeroth order of magnitude is still greater than one, indicating that adiabaticity is not maintained. There are therefore, only three regions in which the approximation holds: at and beyond the rest point, slightly before the transition points where $\frac{d\kappa}{dr_{ob}}\approx 0$ and close to the inner horizon. In the SdS limit, however, there being no inner horizon, $\lim_{r\to 0}(\epsilon^{+},\epsilon^{c})=3$. The control function is only less than one in the proximity of the rest point.
@@
In anti-de Sitter space, the ingoing modes are, in general, adiabatic in only two regimes: first is near the event horizon, at the minima of the effective temperature function, second is at the local maxima before the inner horizon. For an SAdS model, both $\epsilon^{+}$ and $\epsilon^{in}$ converge as they approach the singularity. The asymptotic limit is proportional to the mass parameter, and is less than 0.5 for $M<2$. As the mass parameter increases, the cosmological constant becomes less negative (i.e. increases), and the contraction reduces. As the contraction reduces, adiabaticity is maintained to a greater extent inside the event horizon. It also generally appears to be true that the outgoing modes are more adiabatic than the ingoing modes. This is consistent with the graphs from Fig. 8, according to which the magnitude of the outgoing modes is greater than the ingoing modes.
@@
For the most part of the observer’s trajectory, the adiabatic condition does not hold. One may question the implications for the interpretation of $\kappa$. The form of $\kappa$, as in Eq. (25) and used throughout the paper, is derived directly from the definition given by Eq. (18). It does not rely upon the adiabatic approximation. The observer will still see a nonzero particle production, and therein the Hawking radiation. But the radiation will not have a thermal blackbody spectrum. The graphs in Figs. 9 and 10 show the regime in which the radiation is thermal.